Circular No. 6515

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF SATURN C. Roddier and F. Roddier, Institute for Astronomy (IfA), University

of Hawaii; A. Brahic, Observatoire de Paris; and J. E. Graves, M. J. Northcott and T. Owen, IfA, report: “Images of Saturn’s rings, taken in Aug. 1995 with the University of Hawaii adaptive-optics system mounted on the Canada-France-Hawaii telescope, have now been deconvolved and carefully processed. They show evidence for at least nine additional objects orbiting in the F ring. A good orbital fit (including the effects of the Saturnian J_{2} and J_{4} harmonics) was obtained for all of them with a single distance of 140 500 ± 500 km. The following list includes the three objects already announced on IAUC 6407. As before, the longitudes (uncertainty ±1^{o}) are for the epoch 1995 Aug. 10.5 TT (at Saturn) and are measured from the ascending node of Saturn’s equator on the earth’s J2000.0 equator: S/1995 S 11, longitude 302^{o}, estimated radius 12 km; S/1995 S 9, 317^{o}, 16; S/1995 S 12, 320^{o}, 10; S/1995 S 7 = 1995 S 8, 325^{o}, 20; S/1995 S 13, 330^{o}, 12; S/1995 S 14, 46^{o}, 16; S/1995 S 15, 105^{o}, 12; S/1995 S 16, 114^{o}, 10; S/1995 S 17, 116^{o}, 10; S/1995 S 18, 118^{o}, 10; S/1995 S 19, 120^{o}, 10; S/1995 S 5 = 1995 S 10, 131^{o}, 20. We also find some evidence for S/1995 S 11, S/1995 S 15, S/1995 S 16 and S/1995 S 17 in the HST data (cf. IAUC 6243). S/1995 S 12 is part of the possible S/1995 S 9 arc structure mentioned on IAUC 6407, now resolved into two components; we no longer see evidence for arc structures. The 12 objects listed above cover a total longitude range of 135^{o}. Assuming that objects are uniformly distributed, one can estimate that the F ring contains some 32 of them with radii larger than 10 km.”

COMET C/1995 O1 (HALE-BOPP) H. E. Matthews, National Research Council of Canada and Joint As-

tronomy Centre; D. Jewitt, University of Hawaii; and W. M. Irvine, University of Massachusetts, report the detection in C/1995 O1 of HNC[4-3] at 362.6 GHz using the James Clerk Maxwell Telescope on Dec. 1.0 UT. Measured within a 13′′.5 (29 000 km)-diameter beam, the integrated line intensity was 0.23 ± 0.03 T_{a}∗ K km/s. Using nearly simultaneous observations of HCN[4-3], the HNC/HCN production ratio is estimated as 0.07. This is similar to the value measured earlier in C/1996 B2 (IAUC 6353). Since the two comets were observed at quite different heliocentric distances (2.14 and 1.22 AU), the results suggest that the observed HNC is a parent molecule present in the nuclear ices and not a product of photochemistry in the coma. The HNC/HCN ratio is consistent also with an origin for these species in an interstellar cloud (cf. Irvine et al. 1996, Nature 383, 418).

1996 December 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden